You purchased 100 shares of Facebook, Inc. common stock on margin at $169.25 per share.
Assume the initial margin is 50% and the maintenance margin is 30%. One year later, the stock
price closes at $172.58. If the broker’s call loan rate is 2.00%, what is your return on equity?
Note that Facebook pays no dividends.  
Answer #1

ASSUMING CALL LOAN RATE IS 2% ANNUAL:
Return on Equity=(100*172.58-100*169.25-100*169.25*50%*2%)/(100*169.25*50%)=1.9350%

ASSUMING CALL LOAN RATE IS 2% MONTHLY:
Return on Equity=(100*172.58-100*169.25-100*169.25*50%*2%*12)/(100*169.25*50%)=-20.0650%